OpenStudy (anonymous):

Let f(x) = -5x - 4 and g(x) = 6x - 7. Find f(x) + g(x) A. -11x - 11 C. -11x = 3b D. x - 11

OpenStudy (anonymous):

@agent0smith ?

OpenStudy (anonymous):

x-11? right?

OpenStudy (anonymous):

Given: f(x) = -5x - 4 ----(1 and g(x) = 6x - 7 ----(2 Adding (1 and (2 we find: f(x) + g(x) = -5x - 4 + 6x - 7 i.e. f(x) + g(x) = 6x -5x - 7 -4 i.e. f(x) + g(x) = x - 11 is the required answer... Hence D. x - 11 is the correct option

OpenStudy (anonymous):

Thanks :)